# A Knowledge-Driven Diffusion Policy for End-to-End Autonomous Driving Based on Expert Routing

### v1

@@ Abstract
End-to-end autonomous driving remains constrained by the need to generate multi-modal actions, maintain temporal stability, and generalize across diverse scenarios. Existing methods often collapse multi-modality, struggle with long-horizon consistency, or lack modular adaptability. This paper presents KDP, a knowledge-driven diffusion policy that integrates generative diffusion modeling with a sparse mixture-of-experts routing mechanism. The diffusion component generates temporally coherent and multi-modal action sequences, while the expert routing mechanism activates specialized and reusable experts according to context, enabling modular knowledge composition. Extensive experiments across representative driving scenarios demonstrate that KDP achieves consistently higher success rates, reduced collision risk, and smoother control compared to prevailing paradigms. Ablation studies highlight the effectiveness of sparse expert activation and the Transformer backbone, and activation analyses reveal structured specialization and cross-scenario reuse of experts. These results establish diffusion with expert routing as a scalable and interpretable paradigm for knowledge-driven end-to-end autonomous driving. Validation results and demonstrations are available at Our Project Page.
@@ I Introduction
Autonomous driving has emerged as a central challenge in intelligent transportation and machine learning research, with the promise of improving safety, efficiency, and accessibility in mobility systems[1, 2]. Among the various paradigms, end-to-end (E2E) learning, which directly maps raw sensory observations to driving actions, has gained significant attention for its ability to bypass hand-crafted modular pipelines and exploit the representational power of deep learning models[3, 4].
Despite their promise, existing E2E approaches encounter three fundamental challenges. The first concerns the multi-modality of driving decisions: in many situations, several distinct but equally valid maneuvers can be taken, such as choosing between overtaking and braking. The second relates to temporal stability over long horizons, since even minor prediction errors may accumulate over time and propagate into unsafe trajectories. The third involves continual adaptability, since robust policies must generalize to new environments without catastrophic forgetting[5, 6].
@@
To overcome these limitations, we introduce KDP, a Knowledge-Driven Diffusion Policy for End-to-End Autonomous Driving Based on Expert Routing. The proposed framework unifies the strengths of diffusion models and MoE. The diffusion component ensures robust handling of multi-modal decisions with temporally consistent trajectories, while the MoE routing mechanism organizes experts as abstract knowledge units. Each expert encodes structured knowledge, such as longitudinal regulation, interaction handling, or lateral negotiation, that can be selectively composed depending on scenario demands. In contrast to task-centric designs, this knowledge-driven framework allows the policy to reuse experts across scenarios while enabling new behaviors to emerge through knowledge combinations. Experiments demonstrate that our approach establishes a modular, generative, and knowledge-driven framework for autonomous driving that can addresses the core challenges of stability, adaptability, and multi-modality.
@@
  • Integration of diffusion modeling with expert routing. We combine diffusion policies with MoE, which can maintain long-horizon consistency while adapting to diverse scenarios through modularity and selective reuse of knowledge.
  • Comprehensive empirical validation. We demonstrate the effectiveness of the proposed approach on diverse driving scenarios, achieving superior safety, generalization, and efficiency compared with baselines.
@@ II-C Mixture-of-Experts in Generative AI
MoE architectures have achieved remarkable success in generative artificial intelligence, particularly in large-scale language and vision models[16, 38]. In natural language processing, models such as DeepSeek-v3[39] demonstrate that MoE can scale to trillions of parameters while retaining efficiency, with experts specializing in different linguistic or semantic patterns. In vision and multi-modal learning, MoE has been incorporated into generative tasks such as image synthesis and text-to-image diffusion, enabling different experts to capture distinct modalities or generative styles. These advances highlight MoE as a general mechanism for modularity and diversity in generative modeling.
@@ III-C2 Forward Diffusion of Action Sequences
A key property of the diffusion process is that $a_{t}$ can be expressed in closed form relative to the clean sequence $a_{0}$:

### v2

@@ Abstract
End-to-end autonomous driving remains constrained by the difficulty of producing adaptive, robust, and interpretable decision-making across diverse scenarios. Existing methods often collapse diverse driving behaviors, lack long-horizon consistency, or require task-specific engineering that limits generalization. This paper presents KDP, a knowledge-driven diffusion policy that integrates generative diffusion modeling with a sparse mixture-of-experts routing mechanism. The diffusion component generates temporally coherent action sequences, while the expert routing mechanism activates specialized and reusable experts according to context, enabling modular knowledge composition. Extensive experiments across representative driving scenarios demonstrate that KDP achieves consistently higher success rates, reduced collision risk, and smoother control compared to prevailing paradigms. Ablation studies highlight the effectiveness of sparse expert activation and the Transformer backbone, and activation analyses reveal structured specialization and cross-scenario reuse of experts. These results establish diffusion with expert routing as a scalable and interpretable paradigm for knowledge-driven end-to-end autonomous driving. Validation results and demonstrations are available at Our Project Page.
@@ I Introduction
Autonomous driving has emerged as a central challenge in intelligent transportation and machine learning research, with the promise of improving safety, efficiency, and accessibility in mobility systems [1, 2]. Among the various paradigms, end-to-end (E2E) learning, which directly maps raw sensory observations to driving actions, has gained significant attention for its ability to bypass hand-crafted modular pipelines and exploit the representational power of deep learning models [3].
Despite their promise, current E2E approaches remain constrained in their ability to generalize across the full diversity of real-world driving. Most methods are dominated by data-driven paradigms, whether deterministic policy regression or large-scale imitation learning. While increasing model capacity and dataset size has yielded incremental gains, this trajectory exhibits diminishing returns, which still struggle to extrapolate beyond training distributions, remain brittle in rare or unseen scenarios, and require engineering efforts that scale prohibitively with data volume [4, 5].
@@
To overcome these limitations, we introduce KDP, a Knowledge-Driven Diffusion Policy for end-to-end autonomous driving. KDP integrates diffusion-based generative modeling with sparse expert routing. The diffusion component offers robust distributional expressivity for action generation, while the MoE mechanism organizes experts as abstract driving knowledge units. Each expert encodes structured capabilities—such as longitudinal regulation, interaction handling, or lateral negotiation—that can be selectively composed according to scenario demands. This knowledge-driven framework enables expert reuse across scenarios and supports the emergence of new behaviors through compositionality. Experiments demonstrate that our approach establishes a modular, generative, and knowledge-driven framework for autonomous driving that enhances adaptability and generalization while supporting compositional reuse of knowledge.
@@
  • Integration of diffusion modeling with expert routing. By combining diffusion policies with MoE, KDP achieves expressive action generation together with modular adaptability across diverse scenarios.
  • Dynamic composition and continual adaptability. Sparse activation is integrated into the framework, enabling the model to dynamically compose driving behaviors by combining reusable knowledge units, which facilitates long-term adaptability and transferability across multiple driving environments.
@@ II-C Mixture-of-Experts in Generative AI
MoE architectures have achieved remarkable success in generative artificial intelligence, particularly in large-scale language and vision models [15, 37]. In natural language processing, models such as DeepSeek-v3 and DeepSeek-R1 [38, 39] demonstrate that MoE can scale to trillions of parameters while retaining efficiency, with experts specializing in different linguistic or semantic patterns. In vision and multi-modal learning, MoE has been incorporated into generative tasks such as image synthesis and text-to-image diffusion, enabling different experts to capture distinct modalities or generative styles. These advances highlight MoE as a general mechanism for modularity and diversity in generative modeling.
@@ III-C2 Forward Diffusion of Action Sequences
A key property of the diffusion process is that $a_{t}$ can be expressed in the closed form relative to the clean sequence $a_{0}$: